a ballet perfrmance brought in $61800 on he sale of 3000 tickets. if the tickets sold for $14 and $25 how many of each were sold?
Answer by checkley71(8403)

25x+14(3000-x)=61800
25x+42000-14x=61800
11x=61800-42000
11x=19800
x=19800/11
x=1800 tickes for $25 were sold &
3000-1800=1200 $14 tickets were sold.
proof
25*1800+14*1200=61800
45000+16800=61800
61800=61800